OpenStudy (anonymous):

Write the standard form of the hyperbola that has a center of (0, 0), vertices at (3, 0) and (-3, 0), and slopes of the asymptotes are 4/3 and -4/3.

OpenStudy (anonymous):

see ill tell u the method

OpenStudy (anonymous):

the hyperbola is of the form x^2/a^2 + y^2/b^2 = 1

OpenStudy (anonymous):

a = 3 as the vertices are at 3,0 and -3,0

OpenStudy (anonymous):

if the angle bw asymptotes is t, the eccentricity e = sec (2t)

OpenStudy (anonymous):

and we know b^2/a^2 = e^2 - 1

OpenStudy (anonymous):

so find b frm this..get it?

OpenStudy (anonymous):

b is 72/25

OpenStudy (anonymous):

get it??????

OpenStudy (amistre64):

that part tends to get me; why would 'b' not be 4?

OpenStudy (anonymous):

yeah i got it i think

OpenStudy (anonymous):

what is the actual equation though

OpenStudy (amistre64):

x^2 y^2 --- - ---- = 1 right? 9 b^2

OpenStudy (anonymous):

so there is no b^2?

OpenStudy (amistre64):

im thinking b^2 = 16 .... but i could be wrong

OpenStudy (amistre64):

hims smart about this stuff which makes me doubt my b^2 :)

OpenStudy (anonymous):

yeah ur right amistre...

OpenStudy (amistre64):

yay!! im right

OpenStudy (anonymous):

im lovesick 2 amistre..bound to go wrong smwhere..lol

OpenStudy (amistre64):

lol trees pic is cute...

OpenStudy (anonymous):

y du think i jumped onto her question? dont mind tree..compliments

OpenStudy (anonymous):

:)

OpenStudy (amistre64):

i thought maybe to give her an answer :)

OpenStudy (anonymous):

yes..bt with ulterior motives 2..lol

OpenStudy (anonymous):

lol :))))))))))))))))

OpenStudy (anonymous):

never knew beautiful people could give parentheses life with their smiles..;)

OpenStudy (anonymous):

lol thanks